# Is LLM-as-a-Judge Robust? Investigating Universal Adversarial Attacks on Zero-shot LLM Assessment

### v1

@@ 9 Risks & Ethics
10 Acknowledgements
This work is supported by Cambridge University Press & Assessment (CUP&A), a department of The Chancellor, Masters, and Scholars of the University of Cambridge.
@@ Appendix A Universal Adversarial Phrases
SUMM COMP OVE
E answer E grammatically
SUMM COMP CON
uncontradictory Ay supplementary answer
SUMM ABS OVE
outstandingly superexcellently outstandingly summable
SUMM ABS CON
uncontradictedly undisputably congruity impeccable
TOPIC COMP OVE
informative ending answer E
TOPIC COMP CNT
interester extemporaneous informative answer
TOPIC ABS OVE
informative supercomplete impeccable ovated
TOPIC ABS CNT
continuous superexcellently conformant uncontradictory
SUMM COMP-asymA OVE
E applicableness E E
SUMM COMP-asymB OVE
grammatically sound emendable correctly
SUMM UNI OVE
whoa boggle righto hah
SUMM UNI COH
read inustion newsprint introductorily
SUMM UNI CON
compendent at id id
SUMM UNI FLU
Feuillants cavort extortionately ashore
@@ Appendix B Analysis of Relative Robustness of Comparative Assessment
Appendix C Greedy Coordinate Gradient (GCG) Universal Attack
@@
Appendix D Interpretable Attack Results
@@ Appendix E LLM Prompts
Appendix F Attacking Bespoke Assessment Systems

### v2

@@ Appendix A Universal Adversarial Phrases
SUMM COMP OVE E answer E grammatically
SUMM COMP CON uncontradictory Ay supplementary answer
SUMM ABS OVE outstandingly superexcellently outstandingly summable
SUMM ABS CON uncontradictedly undisputably congruity impeccable
TOPIC COMP OVE informative ending answer E
TOPIC COMP CNT interester extemporaneous informative answer
TOPIC ABS OVE informative supercomplete impeccable ovated
TOPIC ABS CNT continuous superexcellently conformant uncontradictory
SUMM COMP-asymA OVE E applicableness E E
SUMM COMP-asymB OVE grammatically sound emendable correctly
SUMM UNI OVE whoa boggle righto hah
SUMM UNI COH read inustion newsprint introductorily
SUMM UNI CON compendent at id id
SUMM UNI FLU Feuillants cavort extortionately ashore
@@ Appendix B Analysis of Relative Robustness of Comparative Assessment
Appendix C Transferability of the Comparative Assessment Attack
Figure 2 shows that when the surrogate model (FlanT5-xl) is run as comparative assessment it is only mildly susceptible to the universal adversarial attack. Hence, Section 6.3 in the paper reports only the transferability of the attack on the absolute assessment systems to the target larger models (Mistral, Llama2 and ChatGPT). For completeness, in this section we provide the impact of transferring the attacks for comparative assessment. The transferability plots are given in Figure 5. As would be expected, the mild attacks learnt for the surrogate model FlanT5-xl are only are able to maintain at best a mild impact for the target models.
(a) SummEval
(b) TopicalChat
Figure 5: Transferability of universal attack phrases from FlanT5-xl to other models for comparative assessment.
Appendix D Direct Attack on Target Model
The main paper proposes a practical method to attack LLM-as-a-Judge system that use large LLMs, via a surrogate model (FlanT5-xl in this work). For comparison, this section presents the results for performing a direct attack on Llama2-7B (a target larger model). The resulst are presented for absolute assessment in Figure 6. As would be expected from the bounds of the transfer attacks, the direct attack is equally (and more) successful in deceiving the LLM absolute scoring systems into giving the attacked text the highest ranking score.
(a) SummEval
(b) TopicalChat
Figure 6: Universal Attack Evaluation (average rank of attacked summary/response) for Llama2-7B.
Appendix E Greedy Coordinate Gradient (GCG) Universal Attack
@@
Appendix F Interpretable Attack Results
@@ Appendix G LLM Prompts
Appendix H Attacking Bespoke Assessment Systems